# TemplateRL: Structured Template-Guided Reinforcement Learning for LLM Reasoning

### v4

@@ Appendix D Experimental Details
For fair comparison, we benchmark TemplateRL with the following baselines on Qwen2.5-Math-7B: (1) GRPO (Shao et al., 2024), using the same 5.5k training samples; (2) SimpleRL-Zero (Zeng et al., 2025), applying GRPO to 24k samples from GSM8K and MATH; (3) OpenReasoner-Zero (Hu et al., 2025b), employing PPO with 129k samples; (4) PRIME-Zero (Cui et al., 2025a), utilizing implicit process rewards on 150k NuminaMath (Li et al., 2024) queries; (5) Oat-Zero (Liu et al., 2025), introducing Dr.GRPO to mitigate length bias on 8k MATH questions.
@@ E.3 Extension to Multimodal Domains
As shown in Table 2 in the main text, we extend TemplateRL to multimodal domains. We provide more details here. We employ the Geometry3K dataset (Lu et al., 2021), which contains geometry problems with both textual descriptions and visual diagrams, providing a challenging testbed for multimodal reasoning capabilities. We utilize Qwen2.5-VL-3B-Instruct as the base model and follow the EasyR1 pipeline (Zheng et al., 2025) for RL training. Our training configuration maintains consistency with the text-only setting: batch size of 128, generating 16 samples per prompt, with the number of guidance patterns set to 2 (i.e., $|g|=2$). The generation length is configured to 4096 tokens to accommodate detailed multimodal reasoning processes. We evaluate the multimodal extension across five diverse benchmarks spanning different reasoning domains: (1) Multimodal Mathematical reasoning: MathVision Wang et al. (2024a), MathVerse (Zhang et al., 2025), and MathVista (Lu et al., 2023) for multimodal mathematical problem solving; (2) General reasoning: MMMU (Yue et al., 2024) for broad multimodal understanding; and (3) Visual perception: BLINK (Fu et al., 2025) for visual comprehension tasks.
As shown in Table 2, TemplateRL demonstrates consistent improvements over both CoT baseline and vanilla GRPO across most benchmarks. Notably, TemplateRL achieves the best performance on MathVision (29.8 vs. 27.7 CoT baseline, +7.6% improvement) and MathVerse (30.9 vs. 27.3 CoT baseline, +13.2% improvement), indicating that thought-augmented training effectively transfers to multimodal mathematical reasoning scenarios. While performance on MathVista shows some variation, the overall trend suggests that template guidance integration remains beneficial in multimodal settings. These results demonstrate that TemplateRL’s core principle lies in integrating structured reasoning guidance with policy optimization, which generalizes beyond pure text domains to multimodal reasoning tasks. This extension highlights the framework’s potential for broader applications across diverse modalities and reasoning scenarios.
@@ E.4 Extension to Other RL Algorithms
E.5 Ablation Study on the Action Space
@@
$\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$ 83.4 33.3 77.5 64.7
E.6 Sensitivity to Training Data Size
@@
E.7 Sensitivity to Pattern-Matching Metrics
Beyond the Problem Condition Complexity (PCC) metric (Lee and Heyworth, 2000; Salado and Nilchiani, 2014) employed in the main paper for retrieving optimal thought patterns, we explore alternative pattern-matching strategies to assess the robustness of our approach. Specifically, we investigate three additional metrics: (1) Subquestion-count (SC): the number of subquestions an unstructured problem can be decomposed into; (2) Problem semantics (PS): semantic similarity-based matching using sentence embeddings; and (3) Unified ranking: a composite metric that computes PCC, SC, and PS for each incoming question, performs triple relevance ranking across thought templates, and selects the top candidates through weighted averaging.
Action DC ($a_{1}$) SR ($a_{2}$) SA ($a_{3}$) OST ($a_{4}$) CoT ($a_{5}$)
Ratio 31.3% 7.2% 5.7% 30.3% 25.5%
As shown in Table 11, while the unified ranking metric achieves the highest overall performance (66.0 average), the PCC-only approach demonstrates competitive results (64.7 average) with lower computational overhead. The semantic-based matching (PS-only) shows the weakest performance, likely due to the abstract nature of mathematical reasoning patterns that may not be well-captured by surface-level semantic similarity. Given that PCC provides a simple yet effective solution with minimal computational cost, we adopt it as our default pattern-matching metric. Future work could explore more sophisticated weighting schemes for the unified approach to further enhance performance while maintaining efficiency.
@@ E.7 Sensitivity to Pattern-Matching Metrics
PCC-only 83.4 33.3 77.5 64.7
Unified (Equal) 83.9 36.7 77.5 66.0
E.8 Sensitivity to Seed Data Size.
E.9 Statistics on Reasoning Actions
Table 12: Impact of seed data size.
@@ E.9 Statistics on Reasoning Actions
E.10 Statistical Analysis
@@ Appendix G Additional Discussion
Justification for Choosing MCTS.
@@
Task Requirements:  Our task involves constructing complex reasoning chains from a vast combinatorial action space. This setting demands an algorithm that can efficiently search and optimize within such a large space. MCTS is particularly well-suited for this, as it incrementally explores the most promising reasoning trajectories while maintaining flexibility through stochastic sampling.
Limitations of Other Search Methods: We investigated alternatives such as Breadth-First Search (BFS) and Depth-First Search (DFS), and identified several issues:
@@
MCTS Advantages: MCTS offers a superior balance between exploration and exploitation:
@@
Given its ability to handle large search spaces, its robustness to early decision errors, and its structural alignment with sequential reasoning, MCTS emerged as an effective and efficient approach for our reasoning framework.

### v5

@@ Appendix D Experimental Details
For fair comparison, we benchmark TemplateRL with the following baselines on Qwen2.5-Math-7B: (1) GRPO (Shao et al., 2024), using the same 5.5k training samples; (2) SimpleRL-Zero (Zeng et al., 2025), applying GRPO to 24k samples from GSM8K and MATH; (3) OpenReasoner-Zero (Hu et al., 2025a), employing PPO with 129k samples; (4) PRIME-Zero (Cui et al., 2025a), utilizing implicit process rewards on 150k NuminaMath (Li et al., 2024) queries; (5) Oat-Zero (Liu et al., 2025), introducing Dr.GRPO to mitigate length bias on 8k MATH questions.
@@ E.2 Extension to More Models
[…]
Part I: Test-Time Exploration Analysis (w/o RL training)
CoT (pass@16) 80.0 76.8 78.4
[…]
Part II: Post-Training Exploration Analysis (after RL)
GRPO (trained, pass@16) 83.3 80.8 82.0
[…]
Table 8: Exploration Analysis (Pass@16) on highly capable models (Qwen3-32B). Part I examines exploration quality via test-time inference, which mirrors RL rollout sampling. Part II evaluates post-RL performance.
@@ E.3 Extension to Multimodal Domains
As shown in Table 2 in the main text, we extend TemplateRL to multimodal domains. We provide more details and analysis here.
Dataset Construction.
We construct the MMTrain10K dataset based on Geometry3K (Lu et al., 2021), which contains geometry problems with textual descriptions and visual diagrams. To enhance data diversity and problem complexity distribution, we employ Gemini-2.5-Pro to augment the original dataset by systematically adding or removing auxiliary details from problem statements. This process adjusts problem difficulty while preserving geometric relationships and visual diagrams, ultimately expanding the dataset to 10K samples with varied complexity levels (MMTrain10K).
Implementation Setup.
We utilize Qwen2.5-VL-3B-Instruct as the base model and follow the EasyR1 pipeline (Zheng et al., 2025) for RL training. Our training configuration maintains consistency with the text-only setting: batch size 128, 16 samples per question, $|g|=2$ structured guidance templates, and maximum generation length of 4096 tokens to accommodate detailed multimodal reasoning processes.
Evaluation Benchmarks.
We evaluate across five diverse benchmarks: (1) Mathematical reasoning: MathVision Wang et al. (2024a), MathVerse (Zhang et al., 2025), and MathVista (Lu et al., 2023); (2) General understanding: MMMU (Yue et al., 2024); (3) Visual perception: BLINK (Fu et al., 2025).
Results and Analysis.
As shown in Table 2, TemplateRL consistently outperforms both CoT and GRPO across all benchmarks, achieving +8.4% average improvement over GRPO. Notably, TemplateRL excels on mathematical reasoning tasks: MathVision (+12.0% vs. GRPO) and MathVerse (+14.8% vs. GRPO), demonstrating that structured template guidance effectively transfers to multimodal mathematical reasoning. The consistent improvements across diverse tasks (from math to visual perception), validate TemplateRL’s core principle: integrating structured reasoning guidance with policy optimization generalizes beyond text modality to multimodal scenarios.
@@ E.4 Extension to Other RL Algorithms
E.5 Analysis on Highly Capable Models
A natural concern is whether template guidance might restrict the exploration space of larger, more capable models, potentially hindering their inherent reasoning abilities. To address this, we conduct a two-stage analysis on Qwen3-32B with pass@16 evaluation: (1) Exploration Quality Analysis: We compare standard CoT with template-guided test-time inference to assess the quality of exploration spaces before any RL training-this directly reflects the rollout generation quality during RL sampling; (2) Post-Training Performance: We evaluate models after full RL training using CoT, GRPO, and TemplateRL.
As shown in Table 8, template guidance consistently outperforms unrestricted approaches in both stages. In the exploration phase, template-guided sampling achieves 83.3 on AIME25 and 82.8 on GPQA-D, surpassing CoT baseline by 3.3% and 6.0%, respectively. This indicates that templates enhance the diversity and quality of candidate solutions. After RL training, TemplateRL further amplifies this advantage, demonstrating that structured guidance helps models navigate more efficiently toward high-reward regions without limiting their expressiveness. These results confirm that templates enhance rather than restrict effective exploration of solution spaces, even for larger powerful models. Future work will systematically investigate this scaling behavior across model sizes to further validate these preliminary findings.
E.6 Discussion on Template Construction Strategy
For TemplateRL, a natural question is whether the MCTS-based template construction could be replaced with a simpler end-to-end approach where the policy directly generates abstract templates before producing solutions. Actually, our choice of MCTS is motivated by both theoretical considerations and empirical evidence.
Design Rationale.
Our MCTS-based approach is grounded in two key principles:
(1) Quality of Reasoning Guidance. MCTS-extracted templates are derived from verified correct reasoning trajectories obtained through systematic search and validation. The verification process ensures that each template encodes a genuinely effective problem-solving strategy that has been empirically demonstrated to lead to correct solutions. In contrast, end-to-end generation relies on the policy’s current capabilities to produce templates, which may:
  • Generate suboptimal or erroneous templates, especially when the policy is weak or undertrained;
  • Perpetuate existing biases in the model’s reasoning patterns rather than discovering novel strategies;
As demonstrated in Figure 5(b) (Section 4.3), weaker models under standard GRPO training frequently experience training collapse due to sparse positive signals from unstructured exploration. MCTS mitigates this by providing high-quality templates from the outset, enabling stable learning even for less capable models.
(2) Computational Efficiency. While MCTS appears more complex algorithmically, the computational overhead is minimal in practice. Template construction operates on only 500 seed samples—a one-time investment that incurs negligible cost compared to full RL training on 5K samples. This small seed set yields a high-quality template library that immediately stabilizes subsequent RL training across thousands of iterations. In contrast, end-to-end approaches must learn template generation concurrently with solution generation throughout the entire training process, potentially leading to:
  • Unstable training when template quality is poor;
  • Wasted computational budgets on low-quality rollouts;
  • Slower convergence due to dual optimization.
Empirical Validation.
Additionally, we implement and evaluate an end-to-end baseline on Qwen2.5-Math-7B-Base. The end-to-end method follows a two-stage generation process: first, the policy generates an abstract template (e.g., “divide-and-conquer followed by verification”); second, it produces the complete solution conditioned on this template. Both the template and solution are jointly optimized during RL training. Table 9 presents comprehensive results across three benchmarks. We have two findings:
[…]
GRPO (baseline) 16.7 55.0 35.9
End-to-End Generation 23.3 62.5 42.9
[…]
Table 9: Comparison of template construction methods on Qwen2.5-Math-7B-Base. End-to-end generates templates via policy before solving; TemplateRL uses MCTS-verified templates.
@@ E.6 Discussion on Template Construction Strategy
[…]
(1) MCTS Significantly Outperforms End-to-End. TemplateRL with MCTS-based templates achieves 55.4 average performance, substantially higher than end-to-end generation (42.9). The gap is particularly pronounced on AIME (33.3 vs. 23.3, +43.0%), which requires complex multi-step reasoning where high-quality template guidance is most critical. This validates our hypothesis that verified template extraction is more effective than direct policy generation.
(2) End-to-End Still Improves Over Vanilla GRPO. The end-to-end method achieves 42.9 vs. GRPO’s 35.9 (+19.5%), suggesting that incorporating structured template generation is beneficial. However, the quality of templates matters significantly: MCTS provides an additional +29.1% improvement over end-to-end, demonstrating the value of verified guidance.
Future Directions.
While our results demonstrate MCTS’s current superiority, we acknowledge the potential of hybrid approaches. Future work could explore: (1) Warm-start strategies: Initialize with MCTS-constructed templates to ensure stable early training, then gradually transition to end-to-end refinement once the policy has developed sufficient reasoning capabilities; and (2) Curriculum learning: Start with high-confidence MCTS templates for foundational patterns, then allow the policy to discover novel templates for more complex problems.
E.7 Computational Overhead Analysis.
We provide a detailed quantitative breakdown of the additional computational costs introduced by TemplateRL relative to standard GRPO.
Training overhead.
[…]
  • Per-iteration overhead: PCC calculation adds ${\sim}1$s per batch (batch size 128); structured rollout generation takes ${\sim}1.8\times$ the time of unguided sampling due to iterative action prompting.
  • Total training time: TemplateRL requires ${\sim}18$ GPU-hours (500 steps on 5K samples) vs. ${\sim}15$ GPU-hours for GRPO (+20% overhead). However, this yields a +27.4% relative performance gain (Table 1, 55.8 vs. 43.8 average accuracy), achieving 3.1 accuracy points per GPU-hour vs. 2.9 for GRPO–a favorable efficiency tradeoff.
Inference overhead.
  • Template retrieval: ${\sim}0.05$s per question (PCC calculation + distance ranking), negligible compared to generation time.
  • Generation time: Structured rollout takes ${\sim}1.8\times$ unguided generation due to multi-step action prompting, comparable to self-consistency methods that sample multiple paths.
  • Dynamic updates (optional): Add ${\sim}2\%$ additional training time (Table 3) and ${\sim}0.3$s per question at test-time (Table 4), acceptable for accuracy-critical applications.
In summary, the +20% training overhead is well justified by substantial performance gains, and inference overhead remains practical for real-world deployment.
E.8 Ablation Study on the Action Space
E.9 Sensitivity to Training Data Size
@@
[…]
E.10 Sensitivity to Pattern-Matching Metrics
Beyond the Problem Condition Complexity (PCC) metric (Lee and Heyworth, 2000; Salado and Nilchiani, 2014) employed in the main paper for retrieving optimal thought patterns, we explore alternative pattern-matching strategies to assess the robustness of our approach.
Metric Design Rationale.
Our choice of PCC is motivated by its structural focus: unlike semantic embeddings that primarily capture surface-level similarity in problem statements, PCC quantifies the structural complexity of problems based on the number of prior conditions and constraints. This structural feature directly informs the appropriate reasoning strategy—high-PCC problems often require divide-and-conquer approaches to manage multiple interdependent conditions, while simpler low-PCC problems may need only direct reasoning. This explicit complexity-strategy mapping enhances both interpretability (templates are matched based on observable problem structure) and robustness (matching is less sensitive to superficial phrasing variations).
Alternative Metrics Comparison.
To empirically validate PCC’s effectiveness, we investigate three additional pattern-matching strategies:
  • Subquestion-count (SC): The number of subquestions an unstructured problem can be decomposed into, reflecting reasoning granularity;
  • Problem semantics (PS): Semantic similarity-based matching using sentence embeddings (all-MiniLM-L6-v2), representing surface-level textual similarity;
  • Unified ranking (UR): A composite metric that computes PCC, SC, and PS for each incoming question, performs triple relevance ranking across thought templates, and selects top candidates through weighted averaging ($w_{PCC}$=0.5, $w_{SC}$=0.3, $w_{PS}$=0.2).
As shown in Table 12, while the unified ranking metric achieves the highest overall performance (66.0 average), the PCC-only approach demonstrates competitive results (64.7 average) with significantly lower computational overhead. Notably, semantic-based matching (PS-only) shows the weakest performance (60.2 average), which aligns with our hypothesis: mathematical reasoning patterns are inherently structural rather than semantic—two problems may share similar wordings yet require entirely different reasoning strategies, while problems with distinct phrasings may share the same underlying structural complexity. Surface-level semantic similarity fails to capture these critical strategic distinctions, whereas PCC’s structural focus naturally aligns template matching with the actual reasoning demands of problems.
Given that PCC provides a simple, interpretable, and effective solution with minimal computational cost, we adopt it as our default pattern-matching metric. The strong performance of the unified approach (UR) suggests that combining multiple metrics can further enhance matching quality, and future work could explore more sophisticated weighting schemes or learned combination functions to optimize this trade-off between performance and efficiency.
@@ E.10 Sensitivity to Pattern-Matching Metrics
Table 13: Impact of seed data size.
Action DC ($a_{1}$) SR ($a_{2}$) SA ($a_{3}$) OST ($a_{4}$) CoT ($a_{5}$)
Ratio 31.3% 7.2% 5.7% 30.3% 25.5%
E.11 Sensitivity to Seed Data Size.
E.12 Statistics on Reasoning Actions
E.13 Statistical Analysis
@@ Appendix G Additional Discussion
G.1 Justification for Choosing MCTS.
@@
Task Requirements:
Our task involves constructing complex reasoning chains from a vast combinatorial action space. This setting demands an algorithm that can efficiently search and optimize within such a large space. MCTS is particularly well-suited for this, as it incrementally explores the most promising trajectories while maintaining flexibility through stochastic sampling.
Limitations of Other Search Methods:
We investigated alternatives such as Breadth-First Search (BFS) and Depth-First Search (DFS), and identified several issues:
@@
MCTS Advantages:
MCTS offers a superior balance between exploration and exploitation:
@@
Computational Efficiency:
Importantly, MCTS template construction incurs negligible computational overhead in our framework. On 8×A100-80GB GPUs, constructing the template library from 500 seed samples requires about 20 minutes, compared to the $\sim$10-20 hours required for full RL training on 5K samples (500 steps). Moreover, this is a one-time cost: once constructed, the template library can be reused across different model scales, architectures, and even domains (as demonstrated by our cross-domain transfer results). In most scenarios, 500 seed samples are sufficient for building the initial library.
@@
In summary, given its ability to handle large search spaces, its robustness to early decision errors, its structural alignment with sequential reasoning, and its minimal computational cost, MCTS emerged as an effective and efficient approach for our reasoning framework.
G.2 Justification of Key Design Choices
We justify two critical decisions:
(1) Five-action space: The action set {DC, SR, SA, OST, CoT} was designed based on cognitive science literature (Kahneman, 2011) and validated through pilot studies. Testing alternatives on 500-sample validation (diverse task samples) showed: minimal 2-action set (62.4% average accuracy), our 5-action design (71.3%), extended 8-action set (72.4%, only +1.1). The design principle is domain-agnostic meta-strategies rather than domain-specific operations.
(2) PCC over semantic similarity: PCC (Problem Condition Complexity) correctly distinguishes problems by reasoning requirements rather than surface similarity. For example, “find $x$ where $x^{2}+3x-4=0$” (PCC=1) vs. “find $x,y$ where $x+y=5$ and xy=6xy=6” (PCC=2) appear semantically similar but require different strategies (direct CoT vs. DC→CoT).
G.3 Why A Small Set of Generic Patterns Can Generalize?
Our templates encode high-level strategic patterns rather than problem-specific solutions. These templates capture abstract reasoning strategies such as “decompose → solve subproblems → verify” that operate above the surface-level content. While individual problems vary widely in their mathematical details, the underlying reasoning structures often follow recognizable strategic patterns, which is analogous to how humans apply divide-and-conquer or proof-by-contradiction across diverse domains. We believe that not only can individual actions generalize, but their combinations can as well, particularly for complex real-world scenarios where solutions emerge from abstracting more sophisticated action sequences.
G.4 Are Templates Optimized for Test Sets?
No. As shown in Figure 6, templates constructed from mathematical problems generalize effectively to diverse domains: BALROG (+6.1% performance gains for agentic reasoning), GPQA-D (+6.6% gains for graduate-level science), and MMLU-Pro (+7.4% gains for general knowledge). These substantial cross-domain improvements demonstrate that templates are general, not test-set optimized. These templates capture transferable strategic patterns rather than domain-specific solutions.
G.5 How does performance scale with template set size?
In our preliminary experiments, we explored the impact of varying template set sizes and observed a clear pattern: performance improves substantially when increasing template set size below 100 templates, but gains begin to slow beyond this threshold. This suggests an optimal balance point where the template library is sufficiently diverse to cover common reasoning patterns without introducing retrieval complexity or redundancy. Future work could investigate adaptive mechanisms that dynamically adjust the template set based on problem domain or difficulty.
G.6 Robustness to Seed Set and Cross-Domain Generalizability
A potential concern is whether the template library is overly dependent on the quality or domain coverage of the initial 500 seed samples. However, our empirical results demonstrate that the high-level thought patterns abstracted via MCTS possess remarkable task-agnostic properties. As shown in Figure 6, templates derived exclusively from mathematical seeds significantly enhance performance in entirely different domains, such as agentic embodied reasoning (BALROG, +6.1%) and graduate-level science (GPQA-D, +6.6%).
This indicates that TemplateRL captures universal “System 2” strategic logic—including problem decomposition, iterative self-reflection, and constraint analysis—rather than surface-level heuristics. The high-level nature of these patterns mitigates the risk of the model being misled by specific seed biases into suboptimal paths. Therefore, in practical applications, selecting a seed set that covers a wide range of difficulty levels is sufficient to build a robust template library that facilitates effective policy optimization across diverse tasks.
G.7 Distinction between TemplateRL and Related Works
A related line of work, ReMA Wan et al. (2025), learns meta-cognitive strategies through multi-agent reinforcement learning, where agents discover abstract reasoning operators via agent-level reward signals. While ReMA focuses on implicit meta-strategy discovery through agent interactions, TemplateRL constructs explicit, interpretable template libraries via systematic MCTS exploration, enabling immediate human inspection, manual editing, and efficient cross-domain transfer without the multi-agent training overhead.
G.8 Future Work
In future work, we plan to extend TemplateRL along several directions. First, we aim to scale the framework to more complex agentic scenarios Tang et al. (2025); Chen et al. (2025); Wu et al. (2026a); Xu et al. (2026), where dynamic branching and multi-step decision-making demand richer template structures beyond the current five-action space. We also hope to incorporate the entropy mechanism into training to further enhance template diversity Kang et al. (2025). Second, we will explore integrating TemplateRL with heterogeneous model orchestration Wu et al. (2026b), enabling structured template guidance across multi-domain reasoning tasks that require coordinating specialized tools and models. Additionally, we aim to extend templates to more general domains, such as dialogue, retrieval-augmented generation, role-playing, and optimize templates for more diverse objectives Chen et al. (2024); Wu et al. (2025); Liao et al. (2025). When deployed in multi-agent settings, we further plan to investigate how communication topologies influence the propagation of structured templates across agents Shen et al. (2025), and enhance the robustness of template-guided coordination against unknown adversarial conditions Miao et al. (2025). Third, to improve deployment efficiency, we plan to combine TemplateRL with model compression Feng et al. (2025) and inference acceleration techniques Shen et al. (2026), reducing the computational footprint of structured rollout generation while preserving the performance gains demonstrated in this work.